# WaveRoRA: Wavelet Rotary Route Attention for Multivariate Time Series Forecasting

### v1

@@ Abstract
In recent years, Transformer-based models (Transformers) have achieved significant success in multivariate time series forecasting (MTSF). However, previous works focus on extracting features either from the time domain or the frequency domain, which inadequately captures the trends and periodic characteristics. To address this issue, we propose a wavelet learning framework to model complex temporal dependencies of the time series data. The wavelet domain integrates both time and frequency information, allowing for the analysis of local characteristics of signals at different scales. Additionally, the Softmax self-attention mechanism used by Transformers has quadratic complexity, which leads to excessive computational costs when capturing long-term dependencies. Therefore, we propose a novel attention mechanism: Rotary Route Attention (RoRA). Unlike Softmax attention, RoRA utilizes rotary position embeddings to inject relative positional information to sequence tokens and introduces a small number of routing tokens $r$ to aggregate information from the $KV$ matrices and redistribute it to the $Q$ matrix, offering linear complexity. We further propose WaveRoRA, which leverages RoRA to capture inter-series dependencies in the wavelet domain. We conduct extensive experiments on eight real-world datasets. The results indicate that WaveRoRA outperforms existing state-of-the-art models while maintaining lower computational costs.
@@ I Introduction
Multivariate time series (MTS) are prevalent in many real world scenarios, including weather[3], energy management[2], transportation[1] and network flow[9]. Through analyzing historical observations, multivariate time series forecasting (MTSF) aims to extract the underlying patterns of time series data and provides future trends prediction. In the past few years, deep neural networks especially Transformer-based models (Transformers) have achieved great success in MTSF. Transformers rely on the self-attention mechanism[11] to effectively model long-term dependencies in sequences. Models such as Informer[17] and PatchTST[15] make better predictions than previous works that are based on CNNs[5, 4] and RNNs[7]. However, these approaches neglect the correlations between time series, thereby limiting their predictive performance. Some recent works have identified this issue. Models like Sageformer[8] and iTransformer[16] focus on explicitly capturing inter-series dependencies and get superior performance. Nonetheless, the Softmax attention adopted by these models has quadratic computational complexity and may hinder their scalability. Specifically, when modeling intra- or inter-series dependencies, Softmax attention incurs excessive computational costs as the sequence length or the number of variables increases, thus limiting its application in large-scale or high-dimensional scenarios[30, 29, 16].
@@
In addition, existing works mainly focus on analyzing the patterns of MTS data in either time or frequency domains. Features of time domain and frequency domain can reflect different characteristics of the time series. For example, time domain features are conducive to extracting the trends of time series, while frequency domain features reveal the amplitude and phase information of each frequency component of the time series through discrete Fourier transform (DFT)[14]. Compared with time domain features, frequency domain features intuitively reflect the periodicity or seasonality of time series[10]. However, DFT faces the challenges in accurately fitting discontinuous signals, such as abrupt or non-triangular waveforms, which leads to the Gibbs phenomenon[24, 23]. In addition, DFT has limitations in capturing the changes in the frequency domain over time. As shown in Figure 1. Although the three signals exhibit significant differences in the time domain, their frequency spectra are notably similar. Relying exclusively on frequency domain features may result in misleading information and negatively impact the predicting performance of the model. Recent works such as Autoformer[18] and FEDformer[22] introduce frequency-domain features into time-domain representations to facilitate in extracting richer information. However, directly concatenating features of different domains may introduce noise to the model[26], resulting in suboptimal results.
@@
Figure 1: Three signals consisting of the same three functions: $\sin x$, $\sin 3x$ and $\sin 8x$. Signal 1 is defined as $y=\sin x+\sin 3x+\sin 8x$, Signal 2 is evenly divided into 3 segments in the order of $\sin x$, $\sin 3x$ and $\sin 8x$, while signal 3 adjusts the order to $\sin 8x$, $\sin x$ and $\sin 3x$. These signals are different in the time domain but share similar patterns in the frequency domain. The rightmost column shows the wavelet coefficients obtained from a three-level DWT. Compared to the DFT results, the multi-layer wavelet coefficients preserve different periodic characteristics while better revealing the intervals where different periodic patterns are dominant.
@@
Therefore, we propose to extract features in the wavelet domain. Wavelet transform inherits and develops the localization concept of short-time Fourier transform[6], while overcoming its drawbacks such as the fixed window size that does not vary with different frequencies[27]. Wavelet transform extracts the time-frequency characteristics of a time series by scaling and translating a set of localized wavelet basis functions that decay within a short temporal window. For discrete time series, the Discrete Wavelet Transform (DWT)[25] is typically employed, involving a multi-level wavelet decomposition. As illustrated in Fig. 1, the wavelet coefficients produced from the multi-layer decomposition reflect the different periodic characteristics of the time series and accurately reveal the intervals where different periodic patterns are dominant. Our intuition is that extracting features in the wavelet domain allows us to simultaneously leverage the strengths of both time and frequency domains.
@@
We decompose time series into multi-scale high-pass and low-pass components by using DWT. We design a multi-scale wavelet learning framework to facilitate the mutual conversion between wavelet coefficients and unified-size embeddings, aiming to capture temporal patterns within the time series. We then design a novel Rotary Route Attention mechanism (RoRA) and propose WaveRoRA to capture inter-series dependencies in the wavelet domain. We first use rotary position embeddings to inject relative positional information into $Q$ and $K$ matrices. Subsequently, we introduce a routing tokens $r$, which has much fewer elements than the number of variables. The routing tokens aggregate information from the $KV$ matrices and redistribute it to the $Q$ matrix, offering linear complexity. Compared with traditional Softmax attention, RoRA uses the relative positional information to facilitate in identifying the interactions between variables. Meanwhile, the routing tokens effectively alleviate the inherent redundancy between the attention weights and extract critical inter-series dependencies among different variables. Our overall contribution is as follows:
@@
  • We propose to model MTS data in the wavelet domain. We design a multi-scale wavelet learning framework to capture temporal patterns within the time series from high-pass and low-pass components obtained by DWT.
@@
  • We design a novel RoRA mechanism with linear complexity. We use rotary position embeddings to inject relative position information and introduce a small number of routing tokens $r$ to aggregate critical information from the $KV$ matrices and redistribute it to the $Q$ matrix.
@@ III-B Discrete Wavelet Transform
DWT decomposes the signals multiple times to generate wavelets at different resolutions. It can capture the local characteristics of signals in both time and frequency domains, enabling localized analysis of temporal trends and frequency patterns.
@@
Generally, DWT decomposes a given signal $X$ by using a mother function $\psi(t)=a^{-\frac{1}{2}}\psi(\frac{t-b}{a})$ and a father function $\phi(t)=a^{-\frac{1}{2}}\phi(\frac{t-b}{a})$, extracting high-pass and low-pass results respectively at varying resolutions, where $a$ is the scaling factor that controls the dilation of the wavelet and $b$ is the translation factor that shifts the wavelet in time. Specifically, the process of DWT can be defined as Eq 1 and Eq 2, where $j$ indicates the $j$-th decomposition, $yl^{(j)},yh^{(j)}$ represent the low-pass and high-pass results and $yl^{(0)}=X$, $S$ denotes the length of function $\psi$ and $\phi$ and $L^{(j)}=\lfloor\frac{L^{(j-1)}+S-1}{2}\rfloor$ is the length of $yl^{(j)}$. The high-pass and low-pass filters are performed convolution operations with the low-pass results of the former layer, generating temporary outputs $zh$ and $zl$. After the convolution, the results will be downsampled by a factor of 2 according to Nyquist theorem, denoted as $D_{2}$.
$zh^{(j)}(n)$ $=\sum_{i=1}^{S}\psi(i)yl^{(j-1)}(n-i+1)$ (1)
$yh^{(j)}$ $=D_{2}(zh^{(j)})$
$zl^{(j)}(n)$ $=\sum_{i=1}^{S}\phi(i)yl^{(j-1)}(n-i+1)$ (2)
$yl^{(j)}$ $=D_{2}(zl^{(j)})$
@@
There will be $J+1$ wavelet coefficient outputs $\mathcal{C}=\{yh^{(1)},yh^{(2)}...,yh^{(J)},yl^{(J)}\}$ after $J$ layers of decomposition. Given a multivariate input $X\in\mathbb{R}^{L\times M}$, each output contains the wavelet coefficient of each variable, denoted as $yh^{(j)}=[yh^{(j)}_{1},yh^{(j)}_{2},...,yh^{(j)}_{M}]\in\mathbb{R}^{L^{(j)}\times M}$ and $yl^{(j)}=[yl^{(j)}_{1},yl^{(j)}_{2},...,yl^{(j)}_{M}]\in\mathbb{R}^{L^{(j)}\times M}$.
@@
To reconstruct the time series, the Inverse Discrete Wavelet Transform (IDWT) uses the synthesis versions of $\psi$ and $\phi$, denoted as $\psi^{\prime}$ and $\phi^{\prime}$, which are time-reversed forms of $\psi$ and $\phi$. The detailed process of IDWT can be defined in Eq 3:
$\hat{zl}^{(j-1)}=$ $\sum_{i=1}^{S}\psi^{\prime}(i)yh^{(j)}(n-i+1)+$ (3)
$\sum_{i=1}^{S}\phi^{\prime}(i)yl^{(j)}(n-i+1)$
$\hat{yl}^{(j-1)}=$ $U_{2}(\hat{zl}^{(j)})$
@@
where $\hat{zl}$ is the temporary output, $\hat{yl}^{(j)}$ is the low-pass results of the $j$-th layer and $U_{2}$ represents the upsampling operation. The output of $\hat{yl}^{(0)}\in\mathbb{R}^{L\times M}$ is the final reconstructed sequence in the time domain.
Figure 2: The architecture of WaveRoRA. The input MTS data is first stabilized by instance normalization and then transformed to multi-scale wavelet coefficients through $J$-level DWT. Each series of coefficients is passed through a corresponding wave embedding layer to generate uniformed embeddings. These embeddings are then transposed and input into $N$-layer RoRA to capture inter-series dependencies. Subsequently, the output is transposed and fed into a set of wave predictors to obtain the predicted wave coefficients, which are then supplied to the IDWT to generate the final predictions.
@@ III-C Attention Mechanisms
$\mathrm{Attention}(Q,K,V)$ $=\sum_{j=1}^{N}\frac{\mathrm{Sim}(Q_{i},K_{j})}{\sum_{j=1}^{N}\mathrm{Sim}(Q_{i},K_{j})}V_{j},$ (4)
@@
where $\Phi(x)=\mathrm{Elu}(x)+1$ typically. The order of matrix multiplication changes from $Q$-$K$ pairs to $K$-$V$ pairs, leading to $O(Nd^{2})$ complexity. However, linear attention still faces the challenges with excessive computational complexity when dealing with high-dimensional input.
@@ IV-A Overview
As shown in Fig 2, given the input MTS data $\mathbf{X}_{in}\in\mathbb{R}^{L\times M}$, we first apply instance normalization to stabilize the time series. We then apply $J$-level DWT decomposition and get $J$ high-pass coefficients $y_{h}^{(j)}\in\mathbb{R}^{M\times L^{(j)}},j\in\{1,2,...,J\}$ and 1 low-pass coefficients $y_{l}^{(J)}\in\mathbb{R}^{M\times L^{(J)}}$. Due to the downsampling during the multi-layer decomposition of DWT, the lengths of the wavelet coefficients of each layer are inconsistent. Therefore, we introduce a set of wave embedding layers to map the multi-layer coefficients to wavelet embeddings with a uniformed size $D$, denoted as $\mathbb{E}_{j}\in\mathbb{R}^{M\times D}$, where $j=\{1,2,...,J+1\}$. After the projection, we transpose $\mathbb{E}$ into a sequence $\mathbb{S}$ which is composed of individual variables. We use $N$-layer RoRA to capture the inter-series dependencies in the wavelet domain. To simplify the representation, we omit $n\in\{1,2,...,N\}$. For the obtained features of the last layer, we For the obtained features, we flip them again to get $\mathbb{O}$. We aim for each element in $\mathbb{O}$ to correspond to the high-pass and low-pass coefficients obtained after $J$-level wavelet decomposition of the output sequence. Therefore, we again use a set of wave predictor to map them to their corresponding lengths. Finally, we apply IDWT to get the predicted values.

### v2

@@ Abstract
In recent years, Transformer-based models (Transformers) have achieved significant success in multivariate time series forecasting (MTSF). However, previous works focus on extracting features either from the time domain or the frequency domain, which inadequately captures the trends and periodic characteristics. To address this issue, we propose a wavelet learning framework to model complex temporal dependencies of the time series data. The wavelet domain integrates both time and frequency information, allowing for the analysis of local characteristics of signals at different scales. Additionally, the Softmax self-attention mechanism used by Transformers has quadratic complexity, which leads to excessive computational costs when capturing long-term dependencies. Therefore, we propose a novel attention mechanism: Rotary Route Attention (RoRA). Unlike Softmax attention, RoRA utilizes rotary position embeddings to inject relative positional information to sequence tokens and introduces a small number of routing tokens $r$ to aggregate information from the $KV$ matrices and redistribute it to the $Q$ matrix, offering linear complexity. We further propose WaveRoRA, which leverages RoRA to capture inter-series dependencies in the wavelet domain. We conduct extensive experiments on eight real-world datasets. The results indicate that WaveRoRA outperforms existing state-of-the-art models while maintaining lower computational costs. Our code is available at https://github.com/Leopold2333/WaveRoRA.
@@ I Introduction
Multivariate time series (MTS) are prevalent in many real world scenarios, including weather[1], energy management[2], transportation[3] and network flow[4]. Through analyzing historical observations, multivariate time series forecasting (MTSF) aims to extract the underlying patterns of time series data and provides future trends prediction. In the past few years, deep neural networks especially Transformer-based models (Transformers) have achieved great success in MTSF. Transformers rely on the Softmax self-attention (SA) mechanism[5] to effectively model long-term dependencies in sequences. Models such as Informer[6] and PatchTST[7] make better predictions than previous works that are based on CNNs[8, 9] and RNNs[10]. However, these approaches neglect the correlations between time series, thereby limiting their predictive performance. Some recent works have identified this issue. Models like Sageformer[11] and iTransformer[12] focus on explicitly capturing inter-series dependencies and get superior performance. Nonetheless, SA has quadratic computational complexity and may hinder these models’ scalability. Specifically, when modeling intra- or inter-series dependencies, SA incurs excessive computational costs as the sequence length or the number of variables increases, thus limiting its application in large-scale or high-dimensional scenarios[13, 14]. This motivates us to develop a lightweight attention mechanism that simultaneously achieves the expressivity of of SA. In this paper, we design a novel Rotary Route Attention mechanism (RoRA) for MTSF. We first introduce a set of routing tokens $r$ which has much fewer elements than that of the input sequence. The routing tokens aggregate information from the $KV$ matrices and redistribute it to the $Q$ matrix, offering linear complexity. Subsequently, we inject relative positional information through rotary positional embeddings (RoPE)[35] to enhance the feature diversity.
@@
In addition, existing works mainly focus on analyzing the underlying patterns of MTS data in either time or frequency domains. Features of time domain and frequency domain can reflect different characteristics of the time series. For example, time domain features are conducive to extracting the trends of time series, while frequency domain features reveal the amplitude and phase information of each frequency component of the time series through discrete Fourier transform (DFT)[15]. Compared with time domain features, frequency domain features intuitively reflect the periodicity or seasonality of time series[16]. However, DFT faces the challenges in accurately fitting discontinuous signals, such as abrupt or non-triangular waveforms, which leads to the Gibbs phenomenon[17, 30]. In addition, DFT has limitations in capturing the changes in the frequency domain over time. As shown in Figure 1. Although the three signals exhibit significant differences in the time domain, their frequency spectra are notably similar. Relying exclusively on frequency domain features may result in misleading information and negatively impact the predicting performance of the model. Recent works such as Autoformer[18] and FEDformer[19] introduce frequency-domain features into time-domain representations to facilitate in extracting richer information. However, directly concatenating features of different domains may introduce noise to the model[20], resulting in suboptimal results.
@@
Figure 1: Three signals consisting of the same three functions: $\sin x$, $\sin 3x$ and $\sin 8x$ with the sampling rate 20Hz. Signal 1 is defined as $y=\sin x+\sin 3x+\sin 8x$, Signal 2 is evenly divided into 3 segments in the order of $\sin x$, $\sin 3x$ and $\sin 8x$, while signal 3 adjusts the order to $\sin 8x$, $\sin x$ and $\sin 3x$. These signals are different in the time domain but share similar patterns in the frequency domain. The rightmost column shows the wavelet coefficients obtained from a three-level DWT. Compared to the DFT results, the multi-layer wavelet coefficients preserve different periodic characteristics while better revealing the intervals where different periodic patterns are dominant.
@@
Therefore, our intuition is that extracting features in the wavelet domain could be beneficial for more accurate predictions. Wavelet transform inherits and develops the localization concept of short-time Fourier transform[21], while overcoming its drawbacks such as the fixed window size that does not vary with different frequencies[23]. Wavelet transform extracts the time-frequency characteristics of a time series by scaling and translating a set of localized wavelet basis functions that decay within a short temporal window. For discrete time series, the Discrete Wavelet Transform (DWT)[22, 24] is typically employed, involving a multi-level wavelet decomposition. As illustrated in Fig. 1, the wavelet coefficients produced from the multi-layer decomposition reflect the different periodic characteristics of the time series and accurately reveal the intervals where different periodic patterns are dominant. Our intuition is that extracting features in the wavelet domain allows us to simultaneously leverage the strengths of both time and frequency domains.
@@
We decompose time series into multi-scale high-pass and low-pass components by DWT and design a multi-scale wavelet learning framework to facilitate the mutual conversion between wavelet coefficients and unified-size embeddings, aiming to capture temporal patterns within the time series. We then propose WaveRoRA to capture inter-series dependencies in the wavelet domain. Compared with SA, RoRA uses the relative positional information to facilitate in identifying the interactions between variables. Meanwhile, the routing tokens effectively alleviate the inherent redundancy between the attention weights and extract critical inter-series dependencies among different variables. Our contributions can be summarized as follows:
@@
  • We propose modeling MTS data in the wavelet domain. To capture temporal patterns within the time series, we design a multi-scale wavelet learning framework that leverages both high-pass and low-pass components derived from DWT.
@@
  • We design a novel RoRA mechanism with linear complexity, introducing a small number of routing tokens $r$ to aggregate critical information from the $KV$ matrices and redistribute it to the $Q$ matrix. We then use RoPE to inject relative position information.
@@ II Related Work
In this section, we summarize the typical methods of MTSF and provide a simple review of deep models combined with wavelet transform.
@@ III Preliminaries
In this section, we first introduce the concept of multivariate time series forecasting, then summarize the definition of DWT and several attention mechanisms.
@@ III-B Discrete Wavelet Transform
DWT decomposes a given signal $X$ multiple times to generate wavelet coefficients at different resolutions through a high-pass filter $\boldsymbol{h}$ and a low-pass filter $\boldsymbol{g}$. Typically, $\boldsymbol{h}$ and $\boldsymbol{g}$ correspond to a wavelet function $\psi(t)=a^{-\frac{1}{2}}\psi(\frac{t-b}{a})$ and a scaling function $\phi(t)=a^{-\frac{1}{2}}\phi(\frac{t-b}{a})$, which collectively form an orthogonal basis. The parameter $a$ is the scaling factor that controls the dilation of the wavelet while $b$ represents the translation factor that shifts the wavelet in time. Specifically, the process of DWT can be defined as Eq 1 and Eq 2, where $j$ indicates the $j$-th decomposition, $y_{h}^{(j)},y_{l}^{(j)}$ are the high-pass and low-pass components and $y_{l}^{(0)}=X$, $S$ denotes the length of filters and $L^{(j)}=\lfloor\frac{L^{(j-1)}+S-1}{2}\rfloor$ refers to the length of $y_{l}^{(j)}$. $\boldsymbol{h}$ and $\boldsymbol{g}$ are performed convolution operations with the low-pass component of the former layer. After applying the convolutional filters, the frequency range of the signal is compressed to half of its original size. Therefore, the results can be downsampled by a factor of 2 according to the Nyquist theorem. It’s worth noting that the scaling factor $a$ and translation factor $b$ of $\boldsymbol{h}$ and $\boldsymbol{g}$ require no adjustment because (a) the downsampling process simultaneously changes the resolution, allowing for multi-scale analysis; (b) the DWT process inherently encompasses translation as part of the wavelet decomposition.
$y_{h}^{(j)}(n)$ $=\sum_{s=1}^{S}\boldsymbol{h}(s)y_{l}^{(j-1)}(2n-s),~{}n=1,2,...,L^{(j-1)}$ (1)
$y_{l}^{(j)}(n)$ $=\sum_{s=1}^{S}\boldsymbol{g}(s)y_{l}^{(j-1)}(2n-s),~{}n=1,2,...,L^{(j-1)}$ (2)
@@
There will be $J+1$ wavelet coefficient components $\mathcal{C}=\{y_{h}^{(1)},y_{h}^{(2)}...,y_{h}^{(J)},y_{l}^{(J)}\}$ after $J$-level decomposition. Given a multivariate input $X\in\mathbb{R}^{L\times M}$, each output component contains the wavelet coefficient of each variable, denoted as $y_{h}^{(j)}=[y^{(j)}_{h,1},yh^{(j)}_{2},...,yh^{(j)}_{M}]\in\mathbb{R}^{L^{(j)}\times M}$ and $yl^{(j)}=[yl^{(j)}_{1},yl^{(j)}_{2},...,yl^{(j)}_{M}]\in\mathbb{R}^{L^{(j)}\times M}$.
@@
To reconstruct the time series, the Inverse Discrete Wavelet Transform (IDWT) uses the synthesis versions of $\boldsymbol{h}$ and $\boldsymbol{g}$, denoted as $\boldsymbol{h}^{\prime}$ and $\boldsymbol{g}^{\prime}$. The detailed process of IDWT can be defined in Eq 3:
$\hat{y}_{l}^{(j-1)}(n)=$ $\sum_{s=1}^{S}\boldsymbol{h}^{\prime}(s)\tilde{y}_{h}^{(j)}(n-2s)+$ (3)
$\sum_{s=1}^{S}\boldsymbol{g}^{\prime}(s)\tilde{y}_{l}^{(j)}(n-2s),$
$n=$ $1,2,...,2L^{(j)}$
$\tilde{y}_{h}^{(j)}(n)$ =$=\begin{cases}y_{h}^{(j)}(\frac{n}{2}),&\text{if }n\text{ is even}\\ 0,&\text{if }n\text{ is odd}\end{cases}$ (4)
$\tilde{y}_{l}^{(j)}(n)$ =$=\begin{cases}\hat{y}_{l}^{(j)}(\frac{n}{2}),&\text{if }n\text{ is even}\\ 0,&\text{if }n\text{ is odd}\end{cases}$
@@
where $\tilde{y}_{h}^{(j)}$, $\tilde{y}_{l}^{(j)}$ are the upsampled results defined in Eq 4, $\hat{y}_{l}^{(j)}$ is the reconstructed low-pass component of the $j$-th layer and $\hat{y}_{l}^{(J)}=y_{l}^{(J)}$. $\boldsymbol{h}^{\prime}$ and $\boldsymbol{g}^{\prime}$ are performed convolution operations respectively with the high-pass and low-pass components of the current layer. The output of $\hat{yl}^{(0)}\in\mathbb{R}^{L\times M}$ is the final reconstructed sequence in the time domain.
@@ III-C Attention Mechanisms
$\mathrm{Attention}(Q_{i},K_{j},V_{j})$ $=\sum_{j=1}^{N}\frac{\mathrm{Sim}(Q_{i},K_{j})}{\sum_{j=1}^{N}\mathrm{Sim}(Q_{i},K_{j})}V_{j},$ (5)
@@
where $\Phi(x)=\mathrm{Elu}(x)+1$ typically. The order of matrix multiplication changes from $Q$-$K$ pairs to $K$-$V$ pairs, leading to $O(Nd^{2})$ complexity. However, linear attention still faces the challenges with excessive computational complexity when dealing with high-dimensional input data.
Figure 2: The architecture of WaveRoRA. The input MTS data is first stabilized by instance normalization and then transformed to multi-scale wavelet coefficients through $J$-level DWT. Each series of coefficients is passed through a corresponding wave embedding layer to generate uniformed wavelet-wise tokens. These embeddings are then transposed to series-wise tokens and fed into $N$-layer WaveRoRA Encoders to capture inter-series dependencies. Subsequently, the outputs are input to a set of wave predictors to obtain the predicted wavelet coefficients, which are then supplied to the IDWT to generate the final predictions.
@@ IV Methodology
In this section, we overview the model and then detail the operations of each component in the model.
@@ IV-A Overview
As shown in Fig 2, given the input MTS data $\mathbf{X}_{in}\in\mathbb{R}^{L\times M}$, we first apply instance normalization to stabilize the time series. We then apply $J$-level DWT decomposition and get $J$ high-pass coefficients $y_{h}^{(j)}\in\mathbb{R}^{M\times L^{(j)}},j\in\{1,2,...,J\}$ and 1 low-pass coefficients $y_{l}^{(J)}\in\mathbb{R}^{M\times L^{(J)}}$. Due to the downsampling during the multi-layer decomposition of DWT, the lengths of the wavelet coefficients of each layer are inconsistent. Therefore, we introduce a set of wave-embedding layers to map the multi-layer coefficients to wavelet-wise tokens with a uniformed size $D$, denoted as $\mathbb{E}\in\mathbb{R}^{(J+1)\times M\times D}$. After the projection, we transpose $\mathbb{E}$ into series-wise tokens $\mathbb{S}\in\mathbb{R}^{M\times((J+1)\cdot D)}$. We then use $N$-layer WaveRoRA Encoders to capture the inter-series dependencies in the wavelet domain. For the obtained features, we transpose them again to wavelet-wise tokens $\mathbb{O}\in\mathbb{R}^{(J+1)\times M\times D}$. Our objective is that each element in $\mathbb{O}$ approximates the high-pass and low-pass coefficients obtained from $J$-level wavelet decomposition of the ground truth. Therefore, we use a set of wave-predictors to map them to their corresponding lengths. Finally, we apply IDWT to get the predicted values.